OpenStudy (anonymous):

I have a trig final tomorrow and transformations always kick my butt. I understand changing everything to sine and cosine, are there any tips to "see" relationships faster or to generally improve your time and success?

OpenStudy (anonymous):

Or is there a "dummy's guide" of steps?

zepdrix (zepdrix):

What do you mean by transformations? Like given a function \(\large y=2\sin(3x-4)+5\) and being asked to graph it?

OpenStudy (anonymous):

No, more like: Find theta if 0</= theta</= tpi and tan^2 theta - sec theta -1 = 0. I finally figured the answer to that one. Generally solving for theta when you start with garbage so that you can factor or use quadratic in order to graph.

zepdrix (zepdrix):

Ah I see :)

OpenStudy (anonymous):

Graphing it once I have a common value is not a problem it is getting cot, sec, csc sin mises to one form

OpenStudy (anonymous):

"mixes"

OpenStudy (anonymous):

Another transformation example might be: Prove sec (A+B = cos (A+A)/cos^2-sin^2. These types of problems in geral are a pain.

zepdrix (zepdrix):

Well this might seem rather basic, but for recognizing relationships make sure you really really are familiar with the `Square Identities`. sin^2x+cos^2x=1 tan^2x+1=sec^2x cot^2x+1=csc^2x Getting really comfortable with those will definitely help to notice when it's easier to manipulate a square rather than the secant in the problem. Also keep the `Sine Double Angle Formula` in the back of your mind, that one seems to pop up quite a bit. sin(2x)=2 sinx cosx

zepdrix (zepdrix):

Hmmm that's a weird one.. D:

zepdrix (zepdrix):

Is that an actual problem? :D lemme see if i can solve it.

OpenStudy (anonymous):

"I meant cos (A+B)" but that was just an example

zepdrix (zepdrix):

Ah I see :)

OpenStudy (anonymous):

Yes the last one was a real problem, I don't need to do it any more, but I looked it up.

OpenStudy (anonymous):

I guess I just need to work harder on memorizing the basic identities...

zepdrix (zepdrix):

Final is tomorrow? :( hmm.. Trig is my favorite subject.. but it seems to be a tough one for a lot of people. You really have to give yourself a lot of prep time to memorize formulas. There are just soooo many formulas to remember.

OpenStudy (anonymous):

Some of the stuff we are doing in trig the tutor said he didn't see it until calc 3. I lover trig and what you can do with it, I am just too new to it and my placement test put me past algebra which I never had...

zepdrix (zepdrix):

Oh a bit rusty on the algebra? Mmm darn :c